OpenStudy (anonymous):

The product of three numbers in G.P is 1 & the sum of the products of their pairs is 7/2 ; find the numbers

OpenStudy (shubhamsrg):

let the numbers be a,ar,ar^2 now can you form equations ?

OpenStudy (anonymous):

i know what to do first

OpenStudy (anonymous):

yes i can form the 1st one

OpenStudy (anonymous):

what is the 2nd one i cant understand the pair system?

OpenStudy (shubhamsrg):

a(ar) + a(ar^2) + (ar)(ar^2) = 7/2 -> product, in pairs <-

OpenStudy (anonymous):

ok thanks :)

OpenStudy (shubhamsrg):

yw :)

OpenStudy (anonymous):

a*ar*ar^2=1

OpenStudy (shubhamsrg):

yep.

OpenStudy (anonymous):

a^3 r^3=1

OpenStudy (anonymous):

(ar)^3=(1)^3

OpenStudy (shubhamsrg):

You can actually do it quicker, if you assume the 3 terms are a/r , a and ar but a,ar,ar^2 works fine too.

OpenStudy (anonymous):

a/r*a*ar=1

OpenStudy (anonymous):

a=1

OpenStudy (shubhamsrg):

yep, so middle term =1 use second eqn to find r.

OpenStudy (anonymous):

(a/r)(a)+(a/r)(ar)+(a)(ar)=7/2

OpenStudy (shubhamsrg):

a=1,

OpenStudy (shubhamsrg):

make that substitution.

OpenStudy (anonymous):

1+r+r^2=7/2 r

OpenStudy (anonymous):

what now?

OpenStudy (shubhamsrg):

its a quadratic eqn in r, solve for r simply.

OpenStudy (anonymous):

whats the quadratic equation?

OpenStudy (shubhamsrg):

You have never studied about quadratic equations ?

OpenStudy (anonymous):

i meant to say that what quadratic are u getting?

OpenStudy (shubhamsrg):

2 + 2r + 2r^2 = 7r => 2r^2 - 5r +2 =0

OpenStudy (anonymous):

2r^2-(4+1)r+2=0

OpenStudy (shubhamsrg):

yep, go on..

OpenStudy (anonymous):

2r(r-2)-1(r-2)=0

OpenStudy (anonymous):

r=2 & r=1/2

OpenStudy (shubhamsrg):

bingo ! congrats ! ^_^

OpenStudy (anonymous):

so how am i supposed to get the numbers?

OpenStudy (shubhamsrg):

first number = a/r second = a third = ar

OpenStudy (anonymous):

i have to substitute both values of r?

OpenStudy (shubhamsrg):

one by one, you'll get symmetric GPs

OpenStudy (shubhamsrg):

2 of them

OpenStudy (anonymous):

ok!! :)

OpenStudy (anonymous):

thanks

OpenStudy (shubhamsrg):

Glad to help. :)

OpenStudy (anonymous):

1st G.P is 1/2,1,2

OpenStudy (anonymous):

2nd one is 2,1,1/2

OpenStudy (shubhamsrg):

yep.